# thanks: These authors contributed equallythanks: These authors contributed equally

### v2

@@ Abstract
The fractional quantum Hall effect, a paradigmatic topologically ordered state, has been realised in two-dimensional strongly correlated quantum gases and Chern bands of crystals. Here we construct a non-crystalline analogue by coupling quantum wires that are not periodically placed in real-space. Remarkably, the model remains solvable using bosonisation techniques. Due to the non-uniform couplings between the wires, the ground state has different degeneracy compared to the crystalline case. It displays a rich phenomenology of excitations, which can either behave like anyons confined to move in one dimension (lineons), anyons confined to hop between two wires (s-lineons), and anyonic excitations that are free to travel across the system. Both the ground state degeneracy and mutual statistics are directly determined by the real-space positions of the wires. By providing an analytically solvable model of a non-crystalline fractional quantum Hall effect, our work showcases that topological order can display richer phenomenology beyond crystals. More broadly, the non-uniform wire construction we develop can serve as a tool to explore richer many-body phenomenology in non-crystalline systems.
@@ I Introduction
However, there are several examples where strong disorder is not detrimental, and can even be beneficial, for topological order. For instance, the toric code ground state [1], another topologically ordered state, can be defined in any four-fold coordinated graph [1, 12]. It is also possible to define fractional quantum Hall states with trial wave-functions on non-crystalline lattices, for example in fractals [13, 14, 15, 16, 17, 18] and quasicrystals [19].
More strikingly, disorder or non-crystallinity can help to tune into and control the properties of a topologically ordered state. For example, moiré heterostructures, lacking strict lattice periodicity, allow topological bands to flatten by rotating two monolayer materials with respect to one another. By exploiting this freedom, recent experiments have reported the anomalous fractional quantum Hall effect in twisted MoTe2 [20, 21, 22, 23, 24] and rhombohedral multilayer graphene [25, 26, 27, 28], realizing the long-sought prediction of a fractional Hall effect without magnetic fields [29, 30, 31, 32, 33]. Additionally, structural disorder or amorphisation can open a gap in Kitaev’s honeycomb quantum spin-liquid phase leading to a topological, non-abelian chiral spin-liquid phase [34, 35]. This success motivates the question we address in this work.
@@
The coupled-wire construction has been used to construct a rich variety of anisotropic, yet still crystalline, many-body states. These include fractionalised topological phases [107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 104, 135, 136, 137, 138], spin liquids [139, 140, 141, 142, 143, 144, 145, 146, 147, 148] and fractonic states [149, 150, 151, 152, 153, 154]. More recently, the wire construction has also allowed us to analytically describe strongly correlated topological states in twisted moiré heterostructures, as a network of crossing coupled wires [155, 156, 157, 158, 159, 160, 161, 162, 163].
@@
Here we show that this is not the case. By considering coupled wires whose real space configuration is disordered and non-uniform, shown in Fig. 1a, we show that the problem remains solvable via bosonisation. We formalise the constraints on the disordered couplings under which the bosonic fields commute among themselves, while also conserving charge and momentum—thus retaining exact solvability. This problem is equivalent to that of gapping out fractional quantum Hall edge states [9, 164, 165], depicted in Fig. 1b, providing a pathway to realise these novel properties by coupling fractional quantum Hall edges.
@@
We show that the non-crystalline fractional quantum Hall states constructed in this way display features different from their crystalline counterparts. Specifically, we find distinct ground state degeneracy and anyon properties, which depend on the real-space arrangement of the wires. The zoology of anyons includes three types of fractionalised excitations: an excitation constrained to move along the wires (known as lineon in the fractonic language), a composite particle only able to travel up to the next wire and back (which we call a spread-lineon or s-lineon for short), and a composite anyon with complete freedom to travel throughout the system, which we refer to as a C-anyon. Their mutual statistics depends on the real-space arrangement of the wires, that differ in general compared to the crystalline case.
The outline of this paper is as follows: In Section II we lay the foundations for our model, defining the wire construction and its reformulation in terms of bosonised operators. In Section III we show how the degrees of freedom in a single wire may be re-expressed in a basis of fractionalised bosonic operators. Next, we show how couplings may be introduced between adjacent wires that gap out these fractionalised excitations, thus allowing the wires to host a single correlated phase of matter. Importantly, we see that it is possible to couple two adjacent wires even when each hosts differently fractionalised excitations, and discuss the conditions under which this coupling preserves exact solvability and is gapped. Finally, in Section IV we show how this constitutes a general construction of composite and disordered fractional phases, created by gluing many regions with different fractionalisation together. We determine the ground state degeneracy of such a composite system and study the properties of the quasiparticles (both fractonic and non-fractonic) that emerge, determining their charge and braiding statistics.
@@ II The Model
which leads to a parabolic dispersion for the fermions on each wire, with the parabola shifted according to the wire’s position, $E_{j}(k)=\left(k-By_{j}\right)^{2}/2m$. We impose a global Fermi level and corresponding Fermi momentum, $E_{F}=k_{F}/2m$, such that in each wire all momentum states that satisfy $|k_{j}-By_{j}|\leq k_{F}$ are filled, as shown in Fig. 2a.
@@
with $v=\frac{k_{F}}{m}$. Each wire is effectively split into a pair of wires, each with unidirectional transport at a pair of momenta shifted by ±$\pm k_{F}$, as shown in Fig. 2b. Note that Eqs. 5 and 4 should include a shift by the momentum at the Fermi level, $k_{j}^{p}=y_{j}B+(-1)^{p}k_{F}$, where $p\in\{R,L\}$. This shift is removed by redefining $\psi^{a}_{j}(x)\rightarrow e^{ik_{j}^{p}x}\psi^{a}_{j}(x)$.
@@ II.1 Allowed Couplings
where we write ${\psi^{a}_{j}}^{\dagger}$ when $s_{j}^{p}>0$. Since fermionic operators square to zero, these terms must be written using a point-splitting prescription [166]. The coupling Hamiltonian is expressed as a sum over all such coupling terms.
The couplings in Eq. 6 may be represented pictorially as shown in Fig. 3a. It will be convenient to re-express these in terms of two new variables,
@@ II.2 Bosonisation
So far, we have constructed a generic Hamiltonian, represented by a set of linearised dispersive terms, $H^{\textup{Lin}}_{j}$, Eq. 3, and some set of couplings $\mathcal{O}_{\{s_{j}\}}(x)$, Eq. (6). The next step is to write down the bosonic reformulation of this system [167, 168, 166]. We primarily use the conventions of Ref. [166], introducing a set of bosonic fields for left and right movers on each wire, $\phi_{j}^{L/R}$, which satisfy the following commutation relations
@@
Finally, we can write an arbitrary coupling between wires in terms of the bosonic fields, by combining Eqs. 6 and 13
@@ II.2.1 Sum and Difference Fields
In terms of these fields, the linearised Hamiltonian, Eq. 16, takes the form
@@
and each coupling, given by Eq. 17, may be written in terms of the $n_{j}$ and $m_{j}$ variables defined in Eqs. 7 and 8,
@@
Written in terms of bosonic fields, using Eq. 15, this takes the form
[…]
@@ III Fractionalisation
In this section, we shall show how the construction above allows us to produce fractionalised fermionic states. First, we will show in Section III.1 how the left and right-moving fermions on a single wire can be re-expressed in terms of a pair of fractionalised quasiparticle fields. By introducing an appropriate coupling between such fields, we then see how one can stabilise a fractionalised phase, where these quasiparticles make up the low-energy excitations. This discussion largely follows Refs. [102, 103]. Then in Section III.2 we show how it is possible to construct composite materials by connecting wires with different fractionalisation. Crucially, we will see that it is possible to introduce couplings that can gap out left edge modes from one wire with right edge modes from another, where the two wires are fractionalised differently.
@@ III.1 Uniform Fractionalised Phases
These new fields satisfy similar commutation relations to the bare fields, Eq. 12, although they are rescaled by a factor of $u$,
@@
As we shall see in Section IV, these fields correspond to a pair of $1/u$-fractionalised chiral left and right movers.
@@
A diagram of this coupling is shown in Fig. 4a. In terms of bosonic fields, this leads to a term of the form
@@
Comparing to Eqs. 26 and 27, we see that this coupling can be written as
@@
We find that these have almost identical commutation relations to those satisfied by $\theta$ and $\varphi$, see Eqs. 20 and 21, aside from an additional factor of $u$,
@@ III.2 Gluing Fractions Together
Now, let us consider an interface between two sets of coupled wires, where we wish to enforce different fractionalisations ($u_{1}$ and $u_{2}$) in each wire, as shown in Fig. 5. Thus, we wish to couple a $u_{1}$-fractionalised mode, $\tilde{\phi}^{R}_{1}$, from the lower wire with a $u_{2}$-fractionalised mode, $\tilde{\phi}^{L}_{2}$, from the upper wire. This coupling must take an analogous form to Eq. 33, which preserves the exact solubility of the system. With that in mind, let us propose a trial coupling
@@
This coupling determines a pair of operators that play the same role as the $\tilde{\theta}$ and $\tilde{\varphi}$ operators presented in Eqs. 34 and 35,
@@
These should obey an analogous set of commutation relations as Eqs. 36, 37 and 38. In particular, for the coupling to preserve exact solubility, we require that $\tilde{\theta}_{1.5}$ commute with itself at different positions in space. Calculating this explicitly for $\tilde{\theta}$, we find that
@@
for $\tilde{\theta}_{1.5}$ to describe a conserved quantity. Since all variables here must be $\in\mathbb{Z}$. Thus, it is only possible to introduce such a coupling between a pair of fractionalised phases if the ratio between fractions is a square of a rational number. This is consistent with the so-called null-vector criterion proposed by Haldane [9], which is also discussed in detail in [164, 165]. Following this condition, let us write the fractionalisation coefficients $u_{1}$ and $u_{2}$ in terms of a common factor $u$, and our pair of integers $a$ and $b$, according to
@@
We may now write the full commutation relations of the quasiparticle sum and difference operators, analogous to Eqs. 36, 37 and 38,
@@ III.2.1 Charge Conservation
Thus, following the discussion in Section II.1, we see that such a term involves a net addition (or subtraction in the case of the hermitian conjugate) of $(a-b)$ bare fermions to the system. Taken alone, we see that this coupling does not respect charge conservation [164, 165]. However, as is shown in Appendix B, charge conservation may be restored by including an additional non-fractionalised wire which plays the role of a substrate. This exchanges charge with our system, but changes none of the properties of the fractionalised phase, and so we neglect to write explicitly it in what follows.
@@ III.2.2 Momentum Conservation
We may additionally want to check that this coupling satisfies momentum conservation. Comparing with Eq. 11, we see that the condition imposed is
@@ IV A General Construction for Non-Uniform Fractionalisation
We now have all the ingredients necessary to construct a generic composite fractionalised system. That is, given a set of wires, we have shown that it is possible to build couplings that can put each wire in a $1/{u_{j}}$ fractionalised phase iiiOf course, for fermionic systems we require $u_{j}$ odd, however the same can be done starting with bosons to obtain even fractions.. Additionally, we have seen that, provided all fractions take the form $u_{j}=ua_{j}^{2}$, with a common $u$ and all $a_{j}$ coprime, it is possible to introduce a coupling between each wire such that the system forms an overall gapped phase. This construction involves the series of transformations given above, which we compile in Appendix A for ease of use.
We now examine the properties of such a composite construction. In Section IV.1 we determine the degeneracy of the ground state, showing that by mixing fractionalised phases a composite system is able to realise phase with a ground state degeneracy distinct from any of the constituent phases. In Section IV.2 we review the definition of charge carried by every quasiparticle in the uniform Laughlin phase, following [102]. Next, extending to the disordered context, a notion of overall charge cannot be defined (see Section III.2.1), however we define and calculate a quantity analogous to charge, which is conserved. Finally, in Section IV.3 we discuss the braiding of excitations in such a composite fractional material. We explicitly calculate their exchange statistics and show how the system divides into composite fractionalised anyons which are free to move throughout the system, and lineons which can only propagate in the $x$-direction.
Let us start by establishing the most generic configuration, in which we have an array of $N$ wires at positions $y_{j}$. We wish to place each wire in a $1/{u{a_{j}^{2}}}$-fractionalised phase. Thus, we write the degrees of freedom of each wire in terms of left and right-moving fractionalised bosonic fields, $\tilde{\phi}_{j}^{L/R}$, whose commutator includes a factor of $ua_{j}^{2}$, following Eq. 28.
@@
$H=H_{\textup{Kin}}+\int\frac{dx}{2\pi}\sum_{j}\cos\left(2\tilde{\theta}_{j+\frac{1}{2}}(x)\right),$ (56)
where $H_{\textup{Kin}}$ is a quadratic function in $\{\partial_{x}\tilde{\theta}_{j+\frac{1}{2}}\}$ and $\{\partial_{x}\tilde{\varphi}_{j+\frac{1}{2}}\}$. Let us parametrise $H_{\textup{Kin}}$ in terms of a $2N\times 2N$ matrix $M$ as
$H_{Kin}=\int\frac{dx}{2\pi}\begin{pmatrix}\partial_{x}\tilde{\bm{\varphi}}\\ \partial_{x}\tilde{\bm{\theta}}\end{pmatrix}^{T}\begin{pmatrix}M_{\varphi\varphi}&M_{\theta\varphi}\\ M_{\varphi\theta}&M_{\theta\theta}\end{pmatrix}\begin{pmatrix}\partial_{x}\tilde{\bm{\varphi}}\\ \partial_{x}\tilde{\bm{\theta}}\end{pmatrix},$ (57)
@@
Since the coupling is expressed exclusively in terms of the difference field $\tilde{\theta}$, the only condition for the model to remain solvable is that terms involving $\tilde{\varphi}$ (which does not commute with $\tilde{\theta}$) are weak. In the usual bosonisation limit $M_{\varphi\varphi}=M_{\theta\varphi}=M_{\varphi\theta}=0$, so all terms in $H$ commute precisely, and thus we may treat $\tilde{\theta}$ as a classical variable, replacing it with its expectation value. Thus, the system sits in a gapped ground state defined by the lowest energy configuration of $\{\tilde{\theta}_{j+\frac{1}{2}}\}$.
Provided that no terms in $H_{\textup{Kin}}$ involving $\tilde{\varphi}$ are strong enough to close the gap to the $\tilde{\theta}$-ordered ground state, we can effectively treat the system as fluctuating close to this exact, effectively classical, ground state. In what follows, we shall assume that we are working in this limit, the conventional limit where such bosonisation constructions are valid [102, 103, 104].

### v4

@@ Abstract
In non-interacting systems, disorder can drive a trivial phase into a topological one. However little is known how to construct a fractional quantum Hall ground-state, a paradigmatic topologically ordered state, that exists both in crystalline and disordered lattices and is qualitatively different to known topological phases. Here, we propose a general method for building such a phase. This is done by coupling quantum wires placed aperiodically in real-space, where the spatial positioning allows us to tune the inter-wire couplings. We call the emergent phase the Fractonic Fractional Quantum Hall Effect as it displays a rich interplay of fractional quantum Hall physics with fractonic constraints, formed by coupling differently-fractionalised wires into a globally gapped phase. The ground state has an exponential degeneracy in system size, a signature of the emergence of fractons. It displays a rich phenomenology of excitations, which can either behave like anyons confined to move in one dimension (lineons), multiples of which can then hop between two wires (s-lineons) or be free to travel across the system (C-anyons), depending on the multiplicity. Both the ground state degeneracy and mutual statistics are directly determined by the real-space positions of the wires, which can be disordered. Our method provides an analytically solvable pathway to non-crystalline fractional quantum Hall effects and fractonic theories in two-dimensions, examples of which were lacking.
@@ I Introduction
However, there are several examples where strong disorder is not detrimental, and can even be beneficial, for topological order. For instance, the toric code ground state [1], another topologically ordered state, can be defined in any planar graph [1, 12]. It is also possible to define fractional quantum Hall states with trial wave-functions on non-crystalline lattices, for example in fractals [13, 14, 15, 16, 17, 18] and quasicrystals [19].
More strikingly, disorder or non-crystallinity can help to tune into and control the properties of a topologically ordered state. For example, moiré heterostructures, lacking strict lattice periodicity, allow topological bands to flatten by rotating two monolayer materials with respect to one another. By exploiting this freedom, recent experiments have reported the anomalous fractional and integer quantum Hall effect in twisted MoTe2 [20, 21, 22, 23, 24, 25] and rhombohedral multilayer graphene [26, 27, 28, 29], realizing the long-sought prediction of a fractional Hall effect without magnetic fields [30, 31, 32, 33, 34]. Additionally, structural disorder or amorphisation can open a gap in Kitaev’s honeycomb quantum spin-liquid phase leading to a topological, non-abelian chiral spin-liquid phase [35, 36]. This success motivates the question we address in this work.
@@
The coupled-wire construction has been used to construct a rich variety of anisotropic, yet still crystalline, many-body states. These include fractionalised topological phases [108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 105, 136, 137, 138, 139, 140], spin liquids [141, 142, 143, 144, 145, 146, 147, 148, 149, 150] and fractonic states [151, 152, 153, 154, 155, 156]. More recently, the wire construction has also allowed us to analytically describe strongly correlated topological states in twisted moiré heterostructures, as a network of crossing coupled wires [157, 158, 159, 160, 161, 162, 163, 164, 165, 166].
@@
Here we show that this is not the case. By considering coupled wires whose real space configuration is disordered and non-uniform, shown in Fig.˜1a, we show that the problem remains solvable via bosonisation. We formalise the constraints on the disordered couplings under which the bosonic fields commute among themselves—thus retaining exact solvability.By tuning the spacing between wires we show that it is possible to induce hybridisation between differently fractionalised quasiparticles between each pair of adjacent wires, effectively mixing different fractionalisaions to form a global phase with qualitatively new properties. This procedure is possible because one can gap out fractional quantum Hall edge states with different fractionalisations whenever the ratio of fractionalisations is a square number [9, 167, 168], depicted in Fig.˜1b.
@@
We show that the non-crystalline fractional quantum Hall states constructed in this way display features distinct from known crystalline counterparts. Specifically, we find that our system has many features of a gapped fractonic phase. This is remarkable because it is a 2D construction [169, 153, 170] and currently no mechanism is known for producing fractons in 2D [171]. We show that the system has a ground state degeneracy that scales exponentially with system size. Abelian anyonic quasiparticles live at the interface between each pair of wires and carry statistics that depend on the real-space arrangement of their two adjacent wires. Single quasiparticles are partially immobile, where no physical operator can transport a single fracton between wires—and so behave as lineons. We show that despite being immobile alone, bundles of these elementary lineons have less limited mobility, where the multiplicity of the composite defines its transport. We describe two such composite quasiparticles. Lineons may be combined to form an excitation that is able to travel up to the next wire and back (which we call a spread-lineon or s-lineon for short). S-lineons may then be combined to form a composite anyon with complete freedom to travel throughout the system, which we refer to as a C-anyon. The mutual statistics of these excitations depend on the real-space arrangement of the wires, and differ from that of the constituent fractional Hall regions.
The outline of this paper is as follows: In Section˜II we lay the foundations for our model, defining the wire construction and its reformulation in terms of bosonised operators. In Section˜III we show how the degrees of freedom in a single wire may be re-expressed in a basis of fractionalised bosonic operators. Next, we show how couplings may be introduced between adjacent wires that gap out these fractionalised excitations, thus allowing the system of wires to host a single correlated phase of matter. Importantly, we see that it is possible to couple two adjacent wires even when each hosts differently fractionalised excitations, and discuss the conditions under which this coupling preserves exact solvability and is gapped. In Section˜IV we show how this constitutes a general construction of composite and disordered fractional phases, created by gluing many regions with different fractionalisation together. We study the properties of the quasiparticles (both fractonic and non-fractonic) that emerge, determining their charge, reduced mobility and braiding statistics. Finally, in Section˜V we provide a proof of two fractonic properties. We show that the reduced fractonic mobility is completely robust—that is, no well-posed operator can be written down that transports quasibarticles in a way that breaks the fracton constraint. Finally, we prove that the ground state degeneracy of the system scales exponentially with the system size, leaving the stability of the degeneracy to future work.
@@ II The Model
which leads to a parabolic dispersion for the fermions on each wire, with the parabola shifted according to the wire’s position, $E_{j}(k)=\left(k-By_{j}\right)^{2}/2m$. We impose a global Fermi level and corresponding Fermi momentum, $E_{F}=k_{F}/2m$, such that in each wire all momentum states that satisfy $|k_{j}-By_{j}|\leq k_{F}$ are filled, as shown in Fig.˜2a.
@@
with $v=\frac{k_{F}}{m}$. Each wire is effectively split into a pair of wires, each with unidirectional transport at a pair of momenta shifted by $\pm k_{F}$, as shown in Fig.˜2b. Note that Eqs.˜5 and 4 should include a shift by the momentum at the Fermi level, $k_{j}^{p}=y_{j}B+(-1)^{p}k_{F}$, where $p\in\{R,L\}$. This shift is removed by redefining $\psi^{a}_{j}(x)\rightarrow e^{ik_{j}^{p}x}\psi^{a}_{j}(x)$.
@@ II.1 Allowed Couplings
where we write $\psi^{|s_{j}^{p}|}_{j}$$^{\dagger}$ when $s_{j}^{p}>0$. Since fermionic operators square to zero, these terms must be written using a point-splitting prescription [172]. The coupling Hamiltonian is expressed as a sum over all such coupling terms.
The couplings in Eq.˜6 may be represented pictorially as shown in Fig.˜3a. It will be convenient to re-express these in terms of two new variables,
@@ II.2 Bosonisation
So far, we have constructed a generic Hamiltonian, represented by a set of linearised dispersive terms, $H^{\textup{Lin}}_{j}$, Eq.˜3, and some set of couplings $\mathcal{O}_{\{s_{j}\}}(x)$, Eq. (6). The next step is to write down the bosonic reformulation of this system [173, 174, 172]. We primarily use the conventions of Ref. [172], introducing a set of bosonic fields for left and right movers on each wire, $\phi_{j}^{L/R}$, which satisfy the following commutation relations
@@
Finally, we can write an arbitrary coupling between wires in terms of the bosonic fields, by combining Eqs.˜6 and 13
@@ II.2.1 Sum and Difference Fields
In terms of these fields, the linearised Hamiltonian, Eq.˜16, takes the form
@@
and each coupling, given by Eq.˜17, may be written in terms of the $n_{j}$ and $m_{j}$ variables defined in Eqs.˜7 and 8,
@@
Written in terms of bosonic fields, using Eq.˜15, this takes the form
$\begin{aligned} \rho_{j}(x)&=\partial_{x}\phi^{R}(x)-\partial_{x}\phi^{L}(x),\\ &=2\partial_{x}\theta_{j}(x).\end{aligned}$ (25)
@@ III Fractionalisation
In this section, we shall show how the construction above allows us to produce fractionalised fermionic states. First, we will show in Section˜III.1 how the left and right-moving fermions on a single wire can be re-expressed in terms of a pair of fractionalised quasiparticle fields. By introducing an appropriate coupling between such fields, we then see how one can stabilise a fractionalised phase, where these quasiparticles make up the low-energy excitations. This discussion largely follows the standard method of Refs. [103, 104, 106, 105]. Then in Section˜III.2 we show how it is possible to construct composite materials by connecting wires with different fractionalisation. Crucially, we will see that it is possible to introduce couplings that can gap out left edge modes from one wire with right edge modes from another, where the two wires have different fractionalisation.
@@ III.1 Uniform Fractionalised Phases
These new fields satisfy similar commutation relations to the bare fields, Eq.˜12, although they are rescaled by a factor of $u$,
@@
As we shall see in Section˜IV, these fields correspond to a pair of $1/u$-fractionalised chiral left and right movers.
@@
A diagram of this coupling is shown in Fig.˜4a. In terms of bosonic fields, this leads to a term of the form
@@
Comparing to Eqs.˜26 and 27, we see that this coupling can be written as
@@
We find that these have almost identical commutation relations to those satisfied by $\theta$ and $\varphi$, see Eqs.˜20 and 21, aside from an additional factor of $u$,
@@ III.2 Gluing Fractions Together
Now, let us consider an interface between two sets of coupled wires, where we wish to enforce different fractionalisations ($u_{1}$ and $u_{2}$) in each wire, as shown in Fig.˜5. Thus, we wish to couple a $u_{1}$-fractionalised mode, $\tilde{\phi}^{R}_{1}$, from the lower wire with a $u_{2}$-fractionalised mode, $\tilde{\phi}^{L}_{2}$, from the upper wire. This coupling must take an analogous form to Eq.˜33, which preserves the exact solubility of the system. With that in mind, let us propose a trial coupling
@@
This coupling determines a pair of operators that play the same role as the $\tilde{\theta}$ and $\tilde{\varphi}$ operators presented in Eqs.˜34 and 35,
@@
These should obey an analogous set of commutation relations as Eqs.˜36, 37 and 38. In particular, for the coupling to preserve exact solubility, we require that $\tilde{\theta}_{1.5}$ commute with itself at different positions in space. Calculating this explicitly for $\tilde{\theta}$, we find that
@@
we ensure that $\tilde{\theta}_{1.5}$ satisfies the desired commutation relations, and so can have a well-defined expectation value. Since all variables here must be $\in\mathbb{Z}$. Thus, it is only possible to introduce such a coupling between a pair of fractionalised phases if the ratio between fractions is a square of a rational number. This is consistent with the so-called null-vector criterion proposed by Haldane [9], which is also discussed in detail in [167, 168]. Following this condition, let us write the fractionalisation coefficients $u_{1}$ and $u_{2}$ in terms of a common factor $u$, and our pair of integers $a$ and $b$, according to
@@
We may now write the full commutation relations of the quasiparticle sum and difference operators, analogous to Eqs.˜36, 37 and 38,
@@ III.2.1 Charge Conservation
Thus, following the discussion in Section˜II.1, we see that such a term involves a net addition (or subtraction in the case of the hermitian conjugate) of (ab)(a-b) bare fermions to the system. Taken alone, this coupling does not respect charge conservation [167, 168]. Two solutions may be proposed to this problem. Firstly, as is shown in Appendix˜B, charge conservation may be restored by coupling to a superconducting substrate to allow for the exchange of cooper pairs between system and substrate. This may be modelled by including an additional non-fractionalised wire which exchanges charge with our system, but changes none of the properties of the fractionalised phase. Thus, we neglect to include it explicitly in the following analysis. Otherwise, we may also consider, as is discussed in [167], starting with a composite particle which already carries a multiple of an electron’s charge, such that the charge conservation is restored under the couplings shown here. We shall not consider this possibility any further and leave it to future work.
@@ III.2.2 Momentum Conservation
We may additionally want to check that this coupling satisfies momentum conservation. Comparing with Eq.˜11, we see that the condition imposed is
@@ IV A General Construction for Non-Uniform Fractionalisation
We now have all the ingredients necessary to construct a generic composite fractionalised system. That is, given a set of wires, we have shown that it is possible to re-express the degrees of freedom of each wire in a $1/{u_{j}}$ fractionalised basis iiiOf course, for fermionic systems we require $u_{j}$ odd, however the same can be done starting with bosons to obtain even fractions.. Additionally, we have seen that, provided all such fractions are related by a square, taking the form $u_{j}=ua_{j}^{2}$, with a common $u$ and all $a_{j}$ coprime to their adjacent wires, one can to introduce a coupling between each pair of wires such that the system forms an overall gapped phase. This construction involves the series of transformations given above, which we compile in Appendix˜A for ease of use.
We now study the properties of excitations that arise within such a construction, deriving the ‘charge’ carried by these particles, the constraints on their transport and their mutual statistics. In Section˜IV.1 we review the definition of charge carried by every quasiparticle in the uniform Laughlin phase, following [103]. Next, extending to the disordered context, a notion of overall charge cannot be defined (see Section˜III.2.1), however we define and calculate a quantity analogous to charge, which is conserved. Finally, in Section˜IV.2 we discuss the braiding of excitations in such a composite fractional material. We explicitly calculate their exchange statistics and show how the system divides into composite fractionalised anyons which are free to move throughout the system, and lineons which can only propagate in the $x$-direction.
Let us start by establishing the most generic configuration, in which we have an array of $N$ wires at positions $y_{j}$. The couplings to each wire are bind some multiple of a $1/{u{a_{j}^{2}}}$-fractionalised set of left and right movers. Thus, we write the degrees of freedom of each wire in terms of left and right-moving fractionalised bosonic fields, $\tilde{\phi}_{j}^{L/R}$, whose commutator includes a factor of $ua_{j}^{2}$, following Eq.˜28.
@@
$H=H_{\textup{Kin}}+\int\frac{dx}{2\pi}\sum_{j}g\cos\left(2\tilde{\theta}_{j+\frac{1}{2}}(x)\right),$ (56)
where $H_{\textup{Kin}}$ is a quadratic function in $\{\partial_{x}\tilde{\theta}_{j+\frac{1}{2}}\}$, and $\{\partial_{x}\tilde{\varphi}_{j+\frac{1}{2}}\}$, and $g$ is a constant that determines the coupling strength. Let us parametrise $H_{\textup{Kin}}$ in terms of a $2N\times 2N$ matrix $M$ as
$H_{\textup{Kin}}=\int\frac{dx}{2\pi}\begin{pmatrix}\partial_{x}\tilde{\bm{\varphi}}\\ \partial_{x}\tilde{\bm{\theta}}\end{pmatrix}^{T}\begin{pmatrix}M^{\varphi\varphi}&M^{\theta\varphi}\\ M^{\varphi\theta}&M^{\theta\theta}\end{pmatrix}\begin{pmatrix}\partial_{x}\tilde{\bm{\varphi}}\\ \partial_{x}\tilde{\bm{\theta}}\end{pmatrix}$, (57)
@@
Since the coupling is expressed exclusively in terms of the difference field $\tilde{\theta}$, the only condition for our solution to remain valid is that terms involving $\tilde{\varphi}$ (which does not commute with $\tilde{\theta}$) are weak. In the limit $M^{\varphi\varphi}=M^{\theta\varphi}=M^{\varphi\theta}=0$, all terms in $H$ commute precisely, and we may treat $\tilde{\theta}$ as a classical variable, replacing it with its expectation value. Consequently, the system sits in a gapped ground state defined by the lowest energy configuration of these $\{\tilde{\theta}_{j+\frac{1}{2}}\}$ fields.
Provided that no terms in $H_{\textup{Kin}}$ involving $\tilde{\varphi}$ are strong enough to close the gap to the $\tilde{\theta}-$ordered ground state, we can effectively treat the system as fluctuating close to this exact, effectively classical ground state. In what follows, we shall assume that we are working in this limit, the conventional limit where such bosonisation constructions are valid [103, 104, 105]. Furthermore, we must also ensure that our couplings, Eq.˜56, remain relevant under renormalisation group (RG) flow. This depends strongly on the form of $H_{\textup{kin}}$, so we must ensure that it is possible to construct a kinetic term such that the desired couplings are the most relevant possible term that satisfies momentum conservation. This is provided in Appendix˜C.